If equal solutions of 0.1M KMnO4 and 0.1M K2Cr2O7 solutions are allowed to oxidize Fe(2+) to Fe (3+) in acidic medium, then Fe2+ Will be: a) more by KMnO4 b)more by K2Cr2O7 c)equal in both cases

Grade:11

1 Answers

Vikas TU
7 years ago
From MnO4-
MnO4-  +  Fe2+ ---->        Fe3+        +       Mn2+
n1 * 5 = n2 * 1
(0.1V)*5 = n2
n2 = 0.5V
 
From K2Cr2O7,
Cr2O7^2-    +      Fe2+    -------->      Fe3+        +      Cr3+
n0 * 4 = n2 * 1
(0.1V)*4 = n2
n2 =0.4V
Comparing both moles that is n2,
Fe^2+ would be more by KMnO4.